# Clothing agnostic Pre-inpainting Virtual Try-ON

### v1

@@ 5.1 Improved entire pipeline
The improved Masking Pipeline and Generate Skin method-based skin inpainting pipelines do not just complement each function, but also create complementary synergies at the entire pipeline level. In particular, the organic integration of these two modules ef-fectively overcomes the structural limitations of the existing Leffa model, and as a result, natural and high-quality results can be stably produced even in a full-body clothing re-placement scenario.
Since the existing Leffa model specialized in top-centered synthesis, it showed a lim-itation in not being able to properly handle complex clothing structures such as bottoms or dress categories. These limitations result in two key problems: incompleteness of detec-tion of clothing areas and the remaining of existing clothing silhouettes. Therefore, this study introduced a double complementary pipeline structure that considers both the im-provement of masking accuracy and the effect of removing existing clothing.
Masking Pipeline utilizes multi-category information from the Dress Code dataset to perform accurate clothing boundary detection, and based on the results, generates an ap-propriate inpainting mask to remove existing clothing areas. Subsequently, the generate skin pipeline performs a function of blocking existing clothing characteristics included in the input image from interfering with the final output by realistically restoring the skin and short-sleeved silhouette to the removed area. This dual processing structure provides an ideal condition for accurately and consistently reflecting the shape of the reference clothing regardless of the shape of the existing clothing.
For example, in a scenario where a person wearing a long-sleeve shirt wears a sleeveless dress, the Masking Pipeline accurately detects the long-sleeve area from the shoulder to the wrist, and the Skin Inpainting Pipeline naturally restores the skin in the area, thereby accurately reproducing the exposure design of the reference dress. This is not just a partial modification, but a structural improvement of the entire process leading to pre-masking-inpainting-diffusion for whole-body synthesis.
@@ 5.2 Supporting for replacement across different clothing types
This integrated pipeline is designed based on the multi-category structure of the Dress Code dataset, enabling free clothing replacement between two categories: top/bottom, top/dress, and bottom/dress. This function was technically impossible in the top-centered synthesis structure of the existing Leffa model, and it is difficult to implement without the structural improvement proposed in this study. In addition, for complex structural clothing in which the top and bottom are integrated, such as dresses, bottom masking and upper body skin restoration are applied in parallel, enabling natural and flawless synthesis results. This allows users to attempt complex styles that include not only style changes within a single category but also shape transitions between categories, which greatly expands the scope of use of virtual wearing systems.
5.3 Maintaining Model-agnostic properties
On the other hand, another important contribution of this study is that this structural improvement is not dependent on a specific model and maintains the possibility of mod-el-agnostic expansion. Since the proposed improvement methods are not deeply coupled to the internal structure of the Leffa model, but are designed with a modular structure ap-plied to the input image preprocessing and masking generation process, they can be easily integrated into similar diffusion-based models (e.g., HumanDiff, FashionStyleGAN, VI-TONDiffusion, etc.).
In other words, improved masking strategies and generate skin inpainting tech-niques can serve as universal complementary techniques that can dramatically improve the accuracy and style reflection capabilities of existing synthetic models, providing high practical value across virtual wearing techniques.
@@ 6 Conclusion
This study proposed a new pipeline structure, CaP-VTON (Clothing agnostic Pre-inpainting Virtual Try-ON), to overcome the limitations of the existing Leffa-based virtual try-on system. While Leffa showed excellent performance in detailed texture repre-sentation through attention-based flow field learning, it had structural limitations such as bottom detection inaccuracy and existing clothing silhouette remaining. In order to solve these two key problems, this study designed a dual complementary pipeline that inte-grates (1) a multi-category masking pipeline based on Dress Code and (2) a generator skin preprocessing mechanism based on Stable Diffusion.
The masking pipeline has been improved to accurately detect various clothing cate-gories such as top, bottom, and dress, and the skin inpainting pipeline is configured to remove existing clothing areas and naturally restore skin and short-sleeved silhouettes to accurately reflect the style of reference clothing. In particular, the generate skin method contributed to the stable transformation from long-sleeved to short-sleeved or sleeveless clothing by naturally restoring skin areas such as arms, shoulders, and armpits according to the human body’s posture after removing clothing. As a result of the experiment, CAP-VTON had the same similar quantitative index compared to the existing Leffa, and at the same time, achieved an accuracy of 92.5% in the evaluation of consistency based on the accuracy of short-sleeved silhouette synthesis, achieving 15.4%p higher performance than Leffa’s 77.1%. In addition, it was confirmed that CAP-VTON consistently reproduces the sleeve length, pattern, and style of reference clothing, unlike other existing models.
This study has the following contributions to the virtual inspection model. First, the bottom and dress masking functions were successfully implemented to support whole-body synthesis. Second, an input-independent synthesis structure was established by removing the existing clothing interference. Finally, the model-agnostic modular structure secured versatility applicable to various diffusion-based models. This contribu-tion can be used as a foundation technology to build a more precise and practical virtual start-up system in the actual e-commerce environment in the future and has the potential to expand into a variety of fields such as simulation of customized clothes, avatar genera-tion, and clothing design preview. Future research is expected to further expand the prac-ticality and applicability of the CAP-VTON pipeline by further elaborating the processing of complex poses, backgrounds, interactions with non-clothing elements such as accesso-ries and hair, improving inference speed in real-time environments, and integrating it with a dynamic clothing recommendation system based on user feedback.

### v3

@@ 5.1 Improved entire pipeline
The improved masking pipeline and Generate Skin-based skin inpainting pipelines do not just complement each function, but also create complementary synergies at the en-tire pipeline level. In particular, the integration of these two modules effectively alleviated the structural limitations of the existing Leffa model, and produced stable and natural re-sults, especially when long/short-sleeve conversion or partial skin exposure correction was required.
Since the existing Leffa model specialized in top-centered synthesis, it showed a lim-itation in not being able to properly handle complex clothing structures such as bottoms or dress categories. These limitations result in two key problems: the incompleteness of detection of clothing areas and the persistence of existing clothing silhouettes. Therefore, this study introduced a double complementary pipeline structure that considers both the improvement of masking accuracy and the effect of removing existing clothing.
The masking pipeline utilizes multi-category information from the DressCode da-taset to perform accurate clothing boundary detection, and based on the results, generates an appropriate inpainting mask to remove existing clothing areas. Subsequently, the gen-erate skin pipeline performs a function of blocking existing clothing characteristics in-cluded in the input image from interfering with the final output by realistically restoring the skin and short-sleeved silhouette to the removed area. This dual processing structure provides an ideal condition for accurately and consistently reflecting the shape of the ref-erence clothing, regardless of the shape of the existing clothing.
For example, in a scenario wherein a person wearing a long-sleeve shirt wears a sleeveless dress, the masking pipeline accurately detects the long-sleeve area, from the shoulder to the wrist, and the Skin Inpainting Pipeline naturally restores the skin therein, thereby accurately reproducing the exposure design of the reference dress. This is not just a partial modification, but a structural improvement of the entire process, leading to pre-masking–inpainting–diffusion for whole-body synthesis.
@@ 5.2 Supporting for replacement across different clothing types
This integrated pipeline is designed based on the multi-category structure of the DressCode dataset, enabling free clothing replacement between two categories, namely, top/bottom, top/dress, and bottom/dress. This function was technically impossible in the top-centered synthesis structure of the existing Leffa model, and it is difficult to implement without the structural improvement proposed in this study. In addition, for structurally complex clothing in which the top and bottom are integrated, such as dresses, bottom masking and upper body skin restoration are applied in parallel, enabling natural and flawless synthesis results. This allows users to produce complex styles that include not only style changes within a single category, but also shape transitions between categories, which greatly expands the scope of use of virtual wearing systems.
5.3 Clothing-agnostic properties
The existing virtual try-on models have limitations, such as not being able to com-pletely remove the silhouette of the clothes in the input image, or being synthesized into the bottom area. For example, when short-sleeved clothes are synthesized on a person wearing a long-sleeved shirt, the arm outline of the original image remains very small, resulting in unnatural overlapping in the resulting image.
In this study, to address these limitations, the existing clothing silhouette was re-moved using a skin inpainting pipeline based on the DressCode masking pipeline and Generate Skin method, and the natural arm and skin areas suitable for the body structure of the original person were restored. Through this, regardless of what type of clothing the input image features, a clothing-agnostic model that can synthesize new reference cloth-ing while the clothing silhouette of the original image is completely removed has been implemented. This approach provides more realistic and natural results even in the future synthesis of skin-exposing clothing such as short-sleeves and sleeveless clothes. Therefore, the cloning-agnostic design of this study serves as an important basis for enabling stable clothing conversion, even in various exposure forms or pose deformation situations, in the future
@@ 6 Conclusion
This study has proposed a new pipeline structure, CaP-VTON (Clothing Agnostic Pre-Inpainting Virtual Try-On), to overcome the limitations of the existing Leffa-based virtual try-on system. While Leffa showed excellent performance in detailed texture repre-sentation through attention-based flow field learning, it had structural limitations such as bottom detection inaccuracy and the persistence of existing clothing silhouettes. To solve these two key problems, this study designed a dual complementary pipeline that inte-grates (1) a multi-category masking pipeline based on DressCode and (2) a Generate Skin preprocessing mechanism based on Stable Diffusion.
The masking pipeline has been improved to accurately detect various clothing cate-gories such as top, bottom, and dress, and the skin inpainting pipeline is configured to remove existing clothing areas and to naturally restore skin and short-sleeved silhouettes to accurately reflect the style of the reference clothing. In particular, the Generate Skin method contributes to stable transformation from long-sleeved to short-sleeved or sleeve-less clothing by naturally restoring skin areas such as arms, shoulders, and armpits, ac-cording to the body’s posture, after removing clothing. The experiment showed that CaP-VTON had the same quantitative index as the existing Leffa, and at the same time, achieved an accuracy of 92.5% in the evaluation of consistency based on the accuracy of short-sleeved silhouette synthesis, achieving a 15.4% higher performance than Leffa’s 77.1%. In addition, it was confirmed that CaP-VTON consistently reproduces the sleeve length, pattern, and style of reference clothing, unlike other existing models.
This study makes the following contributions to the virtual inspection model. First, the bottom and dress masking functions were successfully implemented to support whole-body synthesis. Second, an input-independent synthesis structure was established by removing the interference of existing clothing. Finally, the model-agnostic modular structure secured a versatility that made it applicable to various diffusion-based models. This model thus represents a foundational technology in building a more precise and practical virtual start-up system in the actual e-commerce environment in the future, and has the potential to expand into a variety of fields such as the simulation of customized clothes, avatar generation, and clothing design previews. Future research is expected to further expand the practicality and applicability of the CaP-VTON pipeline by further elaborating the processing of complex poses, backgrounds, and interactions with non-clothing elements such as accessories and hair; improving inference speed in re-al-time environments; and integrating all this with a dynamic clothing recommendation system based on user feedback.